OpenStudy (thecalchater):

Calculus 5 questions.

OpenStudy (thecalchater):

@welshfella @robtobey

OpenStudy (thecalchater):

@ShadowLegendX

OpenStudy (welshfella):

if y = x^7 x = y^(1/7)

OpenStudy (thecalchater):

right so then what?

OpenStudy (welshfella):

so you need to integrate that between y = 0 and y = 128

OpenStudy (thecalchater):

so int from 0 to 128 x=y^1/7 dy?

OpenStudy (welshfella):

No Sorry I misread the question You need the volume of revolution, not the area so you need pi outside the integral and you also need to square y ^1/7

OpenStudy (welshfella):